# ChartAttack: Testing the Vulnerability of LLMs to Malicious Prompting in Chart Generation

### v1

@@ 5.3 Human-based evaluation results
BM25 anchor-positive 40.56 34.45 78.72
all-mpnet-base-v2 MNR anchor 45.28 42.15 80.85
MNR anchor-positive 46.17 42.54 80.14
GISTE anchor-positive 39.33 39.33 79.43
We conduct a pilot study to evaluate the effectiveness of ChartAttack in misleading humans on chart-based QA. We recruit 12 participants and divide them equally into a control group and an experimental group. Each participant answers 25 chart-related questions. Participants in the control group see correct charts, while participants in the experimental group see misleading charts generated by ChartAttack. Overall, participants in the control group achieve 71.2% accuracy on correct charts, whereas participants in the experimental group score 51% on misleading charts, reflecting an average performance decrease of 20.2 pp. This average performance decrease is similar to the one observed in the MLLM-based evaluation on AttackViz (19.7 pp). These results provide preliminary evidence that LLM-generated misleading charts negatively affect human chart comprehension. We provide a detailed description of the human evaluation in Appendix E.
6 Ablation experiments
6.1 Demonstration selection module
6.2 Misleader-generator module
@@ 7 Conclusions
We present a systematic study of how MLLMs can be prompted to generate misleading charts. We introduce ChartAttack, an automated framework for applying design-level misleaders to chart annotations, and show through extensive experiments that such charts substantially degrade chart QA performance across multiple models and datasets. A complementary human study demonstrates that these misleaders also impair human comprehension. To facilitate further research, we release AttackViz, a dataset of paired clean and misleading charts annotated with misleaders and induced misleading answers to chart-related questions. Our findings expose an underexplored attack surface in multimodal chart generation and highlight the need for robustness beyond data-faithful visualization in MLLM-based systems.
We identify two directions for defense. First, AttackViz can be used to fine-tune MLLMs on both correct and misleading charts, helping models learn how misleaders influence question comprehension and improving overall chart understanding. Second, as a black-box attack, ChartAttack can be mitigated with prompt-level defenses such as misleader detection or system-prompt safeguards.
@@
First, our framework and dataset cover only three chart types and rely on Matplotlib for chart generation. While our filtering process reduces issues such as reduced readability in 3D charts or suppressed tick labels in dual-axis charts, some reading errors may still be caused by these plotting limitations rather than the misleader itself.
@@
Our code is released under the Apache 2.0 license. Our dataset combines annotations from PlotQA (Methani et al., 2020) ($C$ BY 4.0) and ChartQA (Masry et al., 2022) (GPLv3). Because ChartQA is GPLv3, the combined dataset is released under GPLv3.
@@ Appendix A Demonstration selection module: Training dataset creation
Appendix B Misleader-generation module: Oracle experiments
We perform oracle experiments to choose the number of demonstrations for the misleader-generator module of ChartAttack. We report results on the validation split of AttackViz. Similar to the ablation used to select the few-shot strategy, we frame this task as a multi-label classification problem, where a chart JSON annotation–question pair may have one or more misleaders. Table B reports results for one-, three-, and five-shot prompting. Moving from one to three shots yields large performance gains across all chart types, with Macro F1 improving from 0.39-0.52 in the 1-shot setting to 0.55-0.92 in the 3-shot setting. Increasing the number of shots from three to five results in smaller but consistent improvements, with Macro F1 rising by up to 0.05 for horizontal bar charts and by 0.13 for line charts. The five-shot setting achieves the highest Macro F1-score for all chart types, indicating improved balance across misleader categories rather than gains driven by dominant labels. Based on these quantitative improvements, we adopt five demonstrations in our final configuration as a practical trade-off between performance and prompt length.
@@ Appendix C AttackViz corpus
C.1 Cross-domain extension
@@
C.2 AttackViz corpus: statistics
Table C.2 summarizes the statistics of the AttackViz corpus across the train, validation, and test splits. The table reports the number of question-chart pairs for each chart type, along with the distribution of misleaders (misleaders) applied to the charts. A dash (-) indicates that a given misleader is not applicable to the corresponding chart type.
@@ C.2 AttackViz corpus: statistics
Horizontal bar 776 40 147 106 - 470 229 106 174
Vertical bar 788 18 123 139 182 424 198 136 169
Line 461 3 286 37 - - - - 224
@@
Horizontal bar 809 57 133 169 - 476 212 97 174
Vertical bar 812 15 148 123 210 429 213 74 199
Line 425 5 278 0 - - - - 192
@@
Horizontal bar 784 54 127 127 - 477 206 84 151
Vertical bar 787 22 147 165 156 413 218 69 193
Line 436 11 285 - - - - - 218
[…]
Horizontal bar 609 3 53 83 - 195 364 1 80
Vertical bar 1682 15 65 158 93 243 1253 56 116
Line 103 12 49 10 - - - - 52
[…]
Horizontal bar 29 0 4 4 - 14 8 0 1
Vertical bar 105 2 4 4 7 35 61 5 19
Line 6 1 4 0 - - - - 3
[…]
Horizontal bar 32 1 2 5 - 11 16 1 4
Vertical bar 138 0 6 9 6 22 100 8 16
Line 19 10 4 9 - - - - 10
Table 5: Statistics of the AttackViz corpus by chart type and misleading technique. #Q denotes the number of questions. Log scale, Line, Stacked, and Color correspond to Inappropriate use of log scale, Inappropriate use of line, Inappropriate use of stacked, and Ineffective color scheme, respectively.
@@ Appendix D Misleader-generator module: Prompt details
Figure 10 presents the task prompt provided to the MLLM in the Misleader-generator module of ChartAttack. The prompt begins by assigning the model a specific role and providing overall task instructions. It then details a multi-step procedure for generating misleading variations of charts, including vertical bar charts, horizontal bar charts, and line charts. This includes guidance on selecting applicable techniques, modifying chart JSON annotations at different levels of complexity, and reasoning about contextual plausibility. The prompt also defines all allowable misleaders, provides examples of minimal annotation modifications for each, and finally describes the expected output format, including how to produce a plausible but incorrect answer.

### v2

@@ 5.3 Human-based evaluation results
We conduct a pilot study to evaluate the effectiveness of ChartAttack in misleading humans on chart QA. We recruit 12 participants, evenly split into control and experimental groups, and each participant answers 25 chart-related questions. The control group views correct charts, while the experimental group sees misleading charts generated by ChartAttack. Participants in the control group achieve 71.2% accuracy, whereas those in the experimental group achieve 51%, corresponding to a 20.2 pp decrease. This decrease is comparable to the 19.7 pp drop observed in the MLLM-based evaluation on AttackViz. These results provide preliminary evidence that LLM-generated misleading charts negatively affect human chart comprehension. Details of the human evaluation are provided in the Appendix F.
6 Mitigation strategies
Prompt-based guard.
We conduct a preliminary defense experiment by adding a system-level guard instruction to the Misleader-generator. The guard warns about adversarial distortions, forbids perceptual manipulation, and instructs the model to treat misleading demonstrations as attacks that must not be followed. We provide the system-level guard prompt in Appendix G.1. We evaluate attack success using ASR_eff, which counts only structurally valid distortions (non-constant scaling factors and inconsistent dual-axis ranges). Table 2 shows that across three attacker settings, the guard prompt does not reduce the attack success rate on the test set, indicating that simple prompt-level safeguards are insufficient against design-level chart attacks.
Fine-tuned MLLM on AttackViz.
We fine-tune the instruct version of Qwen2.5-VL-3B using QLoRA Dettmers et al. (2023) with 4-bit NF4 quantization on the AttackViz dataset to improve robustness to misleading charts. LoRA adapters are applied to the attention and feed-forward layers. Training is implemented using PEFT Mangrulkar et al. (2022) and TRL von Werra et al. (2020). Full training details are provided in Appendix G.2. We compare the fine-tuned model with its quantized instruct base model. The base model achieves 46.09% accuracy on the AttackViz test set, while the fine-tuned model reaches 54.53% (+8.44 pp). As shown in Table 3, performance improves across all misleaders, with gains between +12.47 and +21.16 pp. Performance on correct charts (None) slightly decreases (-3.44 pp), suggesting a trade-off between robustness to misleading charts and performance on standard charts. This trade-off was also observed by Tonglet et al. (2025a) in the context of inference-time mitigation methods.
@@ 6 Mitigation strategies
Qwen (line, Zero-shot) 0.988 0.988
Qwen (v_bar, Few-shot-5) 0.901 0.901
DeepSeek (h_bar, Few-shot-5) 0.727 0.727
Table 2: Effective attack success rate (ASR_eff) with and without the prompt-based guard defense.
[…]
3D 23.58 37.74 +14.16
Dual axis 36.78 54.02 +17.24
Inappropriate use of line 26.92 48.08 +21.16
Inappropriate use of log scale 24.85 39.88 +15.03
Inappropriate use of stacked 18.54 31.01 +12.47
Ineffective color scheme 27.45 42.48 +15.03
Inverted axis 34.17 53.85 +19.68
Misrepresentation 39.86 58.36 +18.50
None 74.89 71.45 -3.44
Table 3: Accuracy of the base and fine-tuned (SFT) Qwen2.5-VL-3B models on the AttackViz test set across misleaders.
@@ 7 Conclusions
We present a systematic study of how MLLMs can be prompted to generate misleading charts. We introduce ChartAttack, an automated framework for applying design-level misleaders to chart annotations, and show through extensive experiments that such charts substantially degrade chart QA performance across multiple models and datasets. A complementary human study provides preliminary evidence that these misleaders can also impair human comprehension. To facilitate further research, we release AttackViz, a dataset of paired clean and misleading charts annotated with misleaders and induced misleading answers to chart-related questions. Our fine-tuning experiments suggest that this dataset can be used to improve the chart understanding capabilities of MLLMs on misleading charts. Our findings expose an underexplored attack surface in multimodal chart generation and highlight the need for robustness beyond data-faithful visualization in MLLM-based systems.
@@
First, our framework and dataset focus on three chart types. However, our chart type selection accounts for a large share of real-world cases: 64% of the misleading charts in the taxonomy proposed by Lo et al. (2022) and 49% of the misleading visualizations in the Misviz benchmark Tonglet et al. (2025b).
@@
Our code is released under the Apache 2.0 license. Our dataset combines annotations from PlotQA (Methani et al., 2020) ($C$ BY 4.0), ChartQA (Masry et al., 2022) (GPLv3) and ChartX (Xia et al., 2025). Because ChartQA is GPLv3, the combined dataset is released under GPLv3.
@@ Appendix A Demonstration selection module: Training dataset creation
Appendix B ChartAttack: Ablation experiments
B.1 Demonstration selection module
[…]
BM25 anchor-positive 40.56 34.45 78.72
all-mpnet-base-v2 MNR anchor 45.28 42.15 80.85
MNR anchor-positive 46.17 42.54 80.14
GISTE anchor-positive 39.33 39.33 79.43
We perform oracle experiments to choose the number of demonstrations for the misleader-generator module of ChartAttack. We report results on the validation split of AttackViz. Similar to the ablation used to select the few-shot strategy, we frame this task as a multi-label classification problem, where a chart JSON annotation–question pair may have one or more misleaders. Table B.1 reports results for one-, three-, and five-shot prompting. Moving from one to three shots yields large performance gains across all chart types, with Macro F1 improving from 0.39-0.52 in the 1-shot setting to 0.55-0.92 in the 3-shot setting. Increasing the number of shots from three to five results in smaller but consistent improvements, with Macro F1 rising by up to 0.05 for horizontal bar charts and by 0.13 for line charts. The five-shot setting achieves the highest Macro F1-score for all chart types, indicating improved balance across misleader categories rather than gains driven by dominant labels. Based on these quantitative improvements, we adopt five demonstrations in our final configuration as a practical trade-off between performance and prompt length.
@@ Appendix C AttackViz corpus
C.1 Misleader selection
Table 6 shows all the misleaders proposed in the taxonomy of (Lo et al., 2022). Each column corresponds to one of the criteria used to select the final subset of misleaders in this work. The criteria are the following:
  • Correct answer unchanged: The misleader does not affect the correct answer to a question associated with the chart.
  • Violates chart grammar: These misleaders break visualization design principles that may lead to incorrect conclusions about the underlying data.
  • Data unchanged: These misleaders do not modify the underlying data table used to generate the chart; therefore, the correct conclusion can still be reached.
  • Python implementable: The misleader can be implemented in Python using Matplotlib.
  • $5+$ occurrences: These misleaders appear frequently in real-world examples.
  • Previously studied: These misleaders have been previously studied in misleading chart QA (Ge et al., 2023; Bharti et al., 2024) or design-support research (Lo et al., 2023)
[…]
Not data [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Selective data [Uncaptioned image] [Uncaptioned image]
Dubious data [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Non sequitur [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Too few data points [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Discretized continuous variable [Uncaptioned image] [Uncaptioned image]
Missing normalization [Uncaptioned image]
Inappropriate item order
Inappropriate metric [Uncaptioned image] [Uncaptioned image]
Questionable prediction [Uncaptioned image] [Uncaptioned image]
Trend line on random data [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Inappropriate use of accumulation [Uncaptioned image] [Uncaptioned image]
Inappropriate aggregation granularity [Uncaptioned image] [Uncaptioned image]
Two-way normalization [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Truncated axis
Dual axis
Inappropriate axis range
Inverted axis
Log scale
Extended axis [Uncaptioned image] [Uncaptioned image]
Data of different magnitudes [Uncaptioned image] [Uncaptioned image]
Linear scale on exponential data [Uncaptioned image] [Uncaptioned image]
Inappropriate use of line chart
Inappropriate use of pie chart [Uncaptioned image] [Uncaptioned image]
Confusing chart type [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Misusing circular layout [Uncaptioned image] [Uncaptioned image]
Inappropriate use of stacked
Inappropriate use of bar chart [Uncaptioned image] [Uncaptioned image]
Inappropriate use of scatterplot [Uncaptioned image] [Uncaptioned image]
Overusing colors
Indistinguishable colors
Color blind unfriendly [Uncaptioned image] [Uncaptioned image]
Missing title [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Missing axis title [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Missing legend [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Missing value labels [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Missing axis [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Missing axis ticks [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Missing units [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Misrepresentation
Inconsistent tick intervals [Uncaptioned image] [Uncaptioned image]
Inconsistent binning size [Uncaptioned image] [Uncaptioned image]
Changing scale [Uncaptioned image] [Uncaptioned image]
Violating color convention [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Inconsistent grouping [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Inconsistent tick labels [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Inconsistent value labels [Uncaptioned image] [Uncaptioned image]
Cluttering [Uncaptioned image] [Uncaptioned image]
Confusing legend [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Plotting error [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Missing abbreviation [Uncaptioned image] [Uncaptioned image]
Misalignment [Uncaptioned image] [Uncaptioned image]
Plotting out of chart [Uncaptioned image] [Uncaptioned image]
Illegible text [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
3D
Area encoding [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ineffective color scheme
Pictorial area encoding [Uncaptioned image]
Inappropriate use of smoothing [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Distractive value labels [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Map projection distortion [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Inappropriate aspect ratio [Uncaptioned image] [Uncaptioned image]
Sine illusion [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Invalid comparison [Uncaptioned image] [Uncaptioned image]
Correlation not causation [Uncaptioned image] [Uncaptioned image]
Pattern seeking [Uncaptioned image] [Uncaptioned image]
Misleading claim [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Misleading annotation [Uncaptioned image] [Uncaptioned image]
Misleading title [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Misleading value labels [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Hidden distribution [Uncaptioned image] [Uncaptioned image]
Overplotting [Uncaptioned image]
Hidden uncertainty [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Hidden population size [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 6: Properties of the misleaders. [Uncaptioned image] indicates that the corresponding property does not apply. Misleaders satisfying all criteria are highlighted in bold.
As shown in Table 6, 13 out of 74 misleaders satisfy all the considered criteria. Moreover, there is substantial overlap between Overusing colors, Indistinguishable colors, and Ineffective color scheme. As a result, we merge these misleaders into a single category, Ineffective color scheme, resulting in a final set of 11 misleaders.
C.2 Cross-domain extension
@@
Table 8 shows the statistics of the ChartX evaluation set. We consider only bar and line charts based on the criteria described in Section 3 and Appendix C. We obtain the chart JSON annotations by extracting the underlying data from the CSV files provided in the dataset.
[…]
v_bar 1224
line 944
Table 8: Distribution of chart types in ChartX.
C.3 AttackViz corpus: statistics
Table C.3 summarizes the statistics of the AttackViz corpus across the train, validation, and test splits. The table reports the number of question-chart pairs for each chart type, along with the distribution of misleaders applied to the charts. A dash (-) indicates that a given misleader is not applicable to the corresponding chart type.
@@ C.3 AttackViz corpus: statistics
Horizontal bar 776 40 147 106 - 470 229 106 174 100 235 37
Vertical bar 788 18 123 139 182 424 198 136 169 67 180 21
Line 461 3 286 37 - - - - 224 - 53 89
@@
Horizontal bar 809 57 133 169 - 476 212 97 174 96 214 40
Vertical bar 812 15 148 123 210 429 213 74 199 68 186 28
Line 425 5 278 0 - - - - 192 - 49 83
@@
Horizontal bar 784 54 127 127 - 477 206 84 151 88 209 45
Vertical bar 787 22 147 165 156 413 218 69 193 79 196 30
Line 436 11 285 - - - - - 218 - 48 87
Table 9: Statistics of the AttackViz corpus by chart type and misleading technique. #Q denotes the number of questions. Log scale, Line, Stacked, Axis range, Item order, and Color correspond to Inappropriate use of log scale, Inappropriate use of line, Inappropriate use of stacked, Inappropriate axis range, Inappropriate item order, and Ineffective color scheme, respectively.
@@ Appendix D Misleader-generator module: Prompt details
Figure 10 presents the task prompt provided to the MLLM in the Misleader-generator module of ChartAttack. The prompt begins by assigning the model a specific role and providing overall task instructions. It then details a multi-step procedure for generating misleading variations of charts, including vertical bar charts, horizontal bar charts, and line charts. This includes guidance on selecting applicable techniques, modifying chart JSON annotations at different levels of complexity, and reasoning about contextual plausibility. Compatibility between misleaders and chart types is explicitly defined following the taxonomy of (Lo et al., 2022) (see Table 1), and the prompt is therefore chart-type-specific, including only misleaders applicable to the given chart type; the context is further defined by the dataset and retrieved examples, avoiding incompatible or ill-defined manipulations. The prompt also enforces minimal modifications, referring to changes only in the annotation fields strictly required to apply a misleader without altering unrelated data or visual properties, which isolates the effects of the misleader and enables controlled perturbations via a rule-based system. Finally, the prompt describes the expected output format, including how to produce a plausible but incorrect answer, and specifies that misleading answers are validated using a consistency filter (Section 4), where numeric answers must exhibit low variance and textual answers must converge to a majority identical response; the final misleading answer is obtained via averaging or majority vote, under the assumption that consistent incorrect responses across models are induced by the applied misleader.